# Infinity categories with duality and hermitian multiplicative infinite loop space machines

### v3

@@ 1 Introduction
Instead of vector bundles one can consider vector bundles equipped with non-degenerate symmetric or anti-symmetric bilinear forms giving rise to the Grothendieck-Witt group and hermitian $K$-theory. Motivated by direct sum $K$-theory we consider in this article a version of direct sum hermitian $K$-theory. Given a symmetric monoidal category ${\mathcal{C}}$ with duality, e.g. finitely generated projective $R$-modules ${\mathcal{P}}(R)$ over a commutative ring $R$ with the direct sum as symmetric monoidal structure and the usual duality, one can consider the maximal ${\mathbf{E}}_{\infty}$-space in the symmetric monoidal category ${\mathcal{H}}^{\mathrm{nd}}({\mathcal{C}})$ of non-degenerate hermitian objects of ${\mathcal{C}}$ and the spectrum associated to its group completion. This version of hermitian $K$-theory coincides with the common definitions in the case of ${\mathcal{P}}(R)$ (see [26] and [27, Theorem A.1] if $2$ is invertible in $R$ and [17, Theorem A] in general, compare also to [18, 1.6. Theoreme]).
Direct sum $K$-theory is also meaningful in the case of modules over ${\mathbf{E}}_{\infty}$-ring spectra, e.g. the direct sum $K$-theory of finitely generated projective $R$-modules for $R$ a connective ${\mathbf{E}}_{\infty}$-ring spectrum is the connective algebraic $K$-theory of $R$ (see e.g. [20, Theorem 5]). We have a similar behavior in the case of hermitian $K$-theory (see Remark 5.38). In this article we exhibit perfect or finitely generated projective modules over an ${\mathbf{E}}_{\infty}$-ring spectrum as an $\infty$-category with duality together with the relevant symmetric monoidal structures, to which direct sum hermitian $K$-theory can be applied, thereby producing an ${\mathbf{E}}_{\infty}$-ring spectrum.
@@
By Theorem 1.3 the right hand side of injection (1) is the set of equivalence classes of dualities on the $\infty$-category of compact spectra. By Lin’s Theorem [19, Theorem 1.1] there is an isomorphism $\pi_{0}(S^{BC_{2}})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$, where ${\mathbb{Z}}_{2}$ is the uncountable ring of 2-adic numbers. This implies that there are uncountably many inequivalent dualities on the $\infty$-category of compact spectra.
@@
We heavily use symmetric monoidal $\infty$-categories [21, Definition 2.0.0.7.]. We identify symmetric monoidal $\infty$-categories with commutative monoid objects in $\mathsf{Cat}_{\infty}$ via [21, Example 2.4.2.4.]. We denote symmetric monoidal $\infty$-categories in the form $({\mathcal{C}},\otimes)$ and call ${\mathcal{C}}$ the underlying $\infty$-category. We typically drop $\otimes$ from the notation unless there are several symmetric monoidal structures on ${\mathcal{C}}$, which we need to distinguish.
@@
There is a forgetful functor $\mathrm{Calg}({\mathcal{C}})\to{\mathcal{C}}$ that preserves small limits [21, Corollary 3.2.2.5.].
A symmetric monoidal $\infty$-category is cartesian if the tensor unit is final and $\otimes$ is the product. By [21, 2.4.1.5., 2.4.1.7.] every $\infty$-category ${\mathcal{C}}$ that admits finite products carries a unique cartesian symmetric monoidal structure denoted by $({\mathcal{C}},\times)$ and there is a canonical equivalence $\mathrm{Calg}({\mathcal{C}},\times)\simeq\mathrm{Cmon}({\mathcal{C}}).$ For every symmetric monoidal $\infty$-category ${\mathcal{C}}$ and commutative algebra $A$ in ${\mathcal{C}}$ the $\infty$-category ${\mathrm{Mod}}_{A}({\mathcal{C}})$ inherits a symmetric monoidal structure from ${\mathcal{C}}$, the relative tensor product $\otimes_{A}$, such that the free $A$-module functor is symmetric monoidal (see [21, Theorem 4.5.2.1.] or [12, Corollary 7.4.]).
@@
We often use the $\infty$-categorical Grothendieck construction ((un)straightening) [22, Theorem 3.2.0.1.], which classifies functors $\alpha:{\mathrm{S}}\to\mathsf{Cat}_{\infty}$ for any small $\infty$-category ${\mathrm{S}}$ by cocartesian fibrations $\mathrm{X}\to{\mathrm{S}}$ resulting in an equivalence
@@
where the right hand side is the subcategory of cocartesian fibrations to ${\mathrm{S}}$ and functors over ${\mathrm{S}}$ preserving cocartesian lifts. A functor ${\mathrm{S}}\to\mathsf{Cat}_{\infty}$ takes values in spaces if and only if the corresponding cocartesian fibration is a left fibration [22, Proposition 2.4.2.4.]. If ${\mathrm{S}}$ is a space, the cocartesian fibration $\mathrm{X}\to{\mathrm{S}}$ is the functor $\mathrm{colim}(\alpha)\to\mathrm{colim}(*)\simeq{\mathrm{S}}$ induced by the unique map $\alpha\to*$ in $\mathrm{Fun}({\mathrm{S}},\mathsf{Cat}_{\infty})$ by [22, Corollary 3.3.4.3.]. Straightening over $BC_{2}$ gives an equivalence $\mathsf{Cat}_{\infty}[C_{2}]\simeq\mathsf{Cat}_{\infty/BC_{2}}$ that sends an $\infty$-category with $C_{2}$-action ${\mathcal{D}}$ to ${\mathcal{D}}_{hC_{2}}$.
Let ${\mathrm{S}}$ be an $\infty$-category with $C_{2}$-action. A $C_{2}$-equivariant functor $\beta:{\mathrm{S}}\to\mathsf{Cat}_{\infty}$, where $\mathsf{Cat}_{\infty}$ carries the trivial action, is classified by a $C_{2}$-equivariant cocartesian fibration $\mathrm{X}\to{\mathrm{S}}$, refining the non-equivariant (un)straightening construction: the $C_{2}$-equivariant functor $\beta$ corresponds to a functor ${\mathrm{S}}_{hC_{2}}\to\mathsf{Cat}_{\infty}$ classified by a cocartesian fibration $\mathrm{Y}\to{\mathrm{S}}_{hC_{2}}$ (over $BC_{2}$), which by straigtening over $BC_{2}$ is precisely classified by a $C_{2}$-equivariant functor $\mathrm{X}\to{\mathrm{S}}$ whose underlying functor (after forgetting the $C_{2}$-action) is a cocartesian fibration [22, Proposition 2.4.2.11.], which classifies the functor $\beta:{\mathrm{S}}\to{\mathrm{S}}_{hC_{2}}\to\mathsf{Cat}_{\infty}$.
@@ 2.1 $\infty$-categories with duality
The $C_{2}$-action on ${\mathcal{C}}$ is classified via straightening by a functor $\phi:{\mathcal{C}}_{hC_{2}}\to BC_{2}$ by [22, Corollary 3.3.4.3.]. By [22, Corollary 3.3.3.2.] there is a canonical equivalence ${\mathcal{C}}^{hC_{2}}\simeq\mathrm{Fun}_{BC_{2}}(BC_{2},{\mathcal{C}}_{hC_{2}})$ and the forgetful functor ${\mathcal{C}}^{hC_{2}}\to{\mathcal{C}}$ identifies with the functor $\kappa:\mathrm{Fun}_{BC_{2}}(BC_{2},{\mathcal{C}}_{hC_{2}})\to{\mathcal{C}}$ evaluating at the unique object of $BC_{2}$. Hence the object $\widetilde{X\times\tau(X)}\in({\mathcal{C}},\tau)^{hC_{2}}$ with the claimed properties is the right Kan extension relative to $BC_{2}$ of $X:*\to{\mathcal{C}}$ along the functor *$*\to BC_{2}$. By [22, Lemma 4.3.2.13.] this right Kan extension exists and sends the unique object of $BC_{2}$ to $X\times\tau(X)$ if the constant functor *×$*\times_{BC_{2}}*\simeq C_{2}\to{\mathcal{C}}_{hC_{2}}$ with value $X$ admits a $\phi$-limit. By [22, Proposition 4.3.1.9.] this $\phi$-limit is the limit of the functor $C_{2}\to{\mathcal{C}}$ corresponding to the pair $(X,\tau(X))$, which is $X\times\tau(X).$
@@ 2.2 Bimonoidal $\infty$-categories with duality
By [21, Corollary 4.8.1.4, Proposition 4.8.1.17.] there is a closed symmetric monoidal structure on $\mathsf{Pr}^{L}$, where the internal hom of ${\mathcal{C}},{\mathcal{D}}\in\mathsf{Pr}^{L}$ is the $\infty$-category $\mathrm{Fun}^{L}({\mathcal{C}},{\mathcal{D}})$ of left adjoint functors ${\mathcal{C}}\to{\mathcal{D}}$.
@@
An object $\mathrm{X}\in\mathrm{Rig}({\mathcal{C}})$ roughly consists of an object $\mathrm{X}\in{\mathcal{C}}$ equipped with two coherently unital, associative and commutative operations $+,\cdot:\mathrm{X}\times\mathrm{X}\to\mathrm{X}$ such that normal-$\cdot$ coherently distributes over +.+.+ .
@@
The $\infty$-category $\mathsf{Cat}_{\infty}^{\prod}$ admits finite products (in fact all small limits) that are preserved by the inclusion $\mathsf{Cat}_{\infty}^{\prod}\subset\mathsf{Cat}_{\infty}$. Since every finite products preserving functor preserves the final object, the $\infty$-category $\mathsf{Cat}_{\infty}^{\prod}$ admits a zero object. For every ${\mathcal{C}}\in\mathsf{Cat}_{\infty}^{\prod}$ let $\mu_{\mathcal{C}}:{\mathcal{C}}\times{\mathcal{C}}\to{\mathcal{C}}$ the functor assigning the two-fold product. By [21, Proposition 2.4.3.19.] the statement follows from the fact that the functor ${\mathcal{C}}\times 0\to{\mathcal{C}}\times{\mathcal{C}}\xrightarrow{\mu_{\mathcal{C}}}{\mathcal{C}}$ is the identity, the functor $({\mathcal{C}}\times{\mathcal{C}})\times({\mathcal{D}}\times{\mathcal{D}})\simeq({\mathcal{C}}\times{\mathcal{D}})\times({\mathcal{C}}\times{\mathcal{D}})\xrightarrow{\mu_{{\mathcal{C}}\times{\mathcal{D}}}}{\mathcal{C}}\times{\mathcal{D}}$ is the functor $\mu_{\mathcal{C}}\times\mu_{\mathcal{D}}$ and for every map $\mathsf{f}:{\mathcal{C}}\to{\mathcal{D}}$ in $\mathsf{Cat}_{\infty}^{\prod}$ we have $\mathsf{f}\circ\mu_{\mathcal{C}}=\mu_{\mathcal{D}}\times(\mathsf{f}\times\mathsf{f}).$
@@
This follows since $\mathsf{Cat}_{\infty}^{\mathrm{preadd}}\subset\mathsf{Cat}_{\infty}^{\prod}$ is closed under finite products. ∎
@@
By [21, Corollary 4.8.1.4] the $\infty$-category $\mathsf{Cat}_{\infty}^{\prod}$ carries a presentably symmetric monoidal structure such that the inclusion $\mathsf{Cat}_{\infty}^{\prod}\subset\mathsf{Cat}_{\infty}$ admits a symmetric monoidal left adjoint. Applying the lax symmetric monoidal localization functor $\mathrm{Cmon}:\mathsf{Pr}^{L}\to\mathsf{Pr}^{L}$ we find that the embedding (2) admits a symmetric monoidal left adjoint and so refines to a lax symmetric monoidal embedding
@@
Let $\mathsf{Cat}_{\infty}^{\coprod}\subset\mathsf{Cat}_{\infty}$ be the subcategory of small $\infty$-categories having finite coproducts and functors preserving finite coproducts. Taking the opposite $\infty$-category restricts to an equivalence $\mathsf{Cat}_{\infty}^{\prod}\simeq\mathsf{Cat}_{\infty}^{\coprod}$, which further restricts to an autoequivalence of $\mathsf{Cat}_{\infty}^{\mathrm{preadd}}.$ Consequently, it is enough to prove that the embedding $\mathsf{Cat}_{\infty}^{\mathrm{preadd}}\subset\mathsf{Cat}_{\infty}^{\coprod}$ admits a symmetric monoidal left adjoint, which is notationally more convenient. For any small $\infty$-category ${\mathcal{C}}$ that admits finite coproducts let ${\mathcal{C}}^{\prime}$ be the essential image of the composition ${\mathcal{C}}\to\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})\to\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))\simeq\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$ of the Yoneda-embedding and the free functor. Since the Yoneda-embedding ${\mathcal{C}}\to\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ preserves finite coproducts [22, Proposition 5.5.8.10.], the full subcategory ${\mathcal{C}}^{\prime}$ is closed in the preadditive $\infty$-category $\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))$ under finite coproducts and therefore is itself preadditive. We will prove that the induced functor ${\mathcal{C}}\to{\mathcal{C}}^{\prime}$ induces an equivalence $\kappa_{\mathcal{D}}:\mathrm{Fun}^{\coprod}({\mathcal{C}}^{\prime},{\mathcal{D}})\to\mathrm{Fun}^{\coprod}({\mathcal{C}},{\mathcal{D}}).$ The functor $\kappa_{\mathcal{D}}$ is the pullback of the functor $\kappa_{\mathrm{Fun}^{\prod}({\mathcal{D}}^{\mathrm{op}},{\mathcal{S}})}$ along the Yoneda-embedding ${\mathcal{D}}\subset\mathrm{Fun}^{\prod}({\mathcal{D}}^{\mathrm{op}},{\mathcal{S}})\simeq\mathrm{Fun}({\mathcal{D}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$. Since $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ is preadditive and presentable [22, Proposition 5.5.8.10.], we can therefore assume that ${\mathcal{D}}$ is a presentable preadditive $\infty$-category when proving that $\kappa_{\mathcal{D}}$ is an equivalence.
@@
The $\infty$-category $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ is generated by ${\mathcal{C}}$ under small sifted colimits [22, Lemma 5.5.8.14.]. Since $\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))$ is generated by $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ under small sifted colimits, $\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))$ is generated by ${\mathcal{C}}^{\prime}$ under small sifted colimits. Every object of ${\mathcal{C}}$ is compact and projective in $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ by [22, Proposition 5.5.8.10.4.] so that every object of ${\mathcal{C}}^{\prime}$ is compact and projective in $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$. By [22, Proposition 5.5.8.22.] this guarantees that the embedding ${\mathcal{C}}^{\prime}\subset\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$ induces an equivalence $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\prime\mathrm{op}},{\mathcal{S}})\simeq\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}})).$
@@
For every small $\infty$-category ${\mathcal{B}}$ that admits finite coproducts the induced functor $\mathrm{Fun}^{\mathrm{L}}(\mathrm{Fun}^{\prod}({\mathcal{B}}^{\mathrm{op}},{\mathcal{S}}),{\mathcal{D}})\to\mathrm{Fun}^{\coprod}({\mathcal{B}},{\mathcal{D}})$ is an equivalence [22, Proposition 5.5.8.15.]. So in the following commutative square the vertical functors are equivalences:
@@
whose fully faithful right adjoint identifies preadditive symmetric monoidal $\infty$-categories with duality $({\mathcal{C}},\otimes,D)$ with preadditive bimonoidal $\infty$-categories with duality $({\mathcal{C}},\otimes,\boxtimes,D)$, where =normal-$\boxtimes$=$\oplus$ = ⊕ is the sum.

### v4

@@ 1 Introduction
Instead of vector bundles one can consider vector bundles equipped with non-degenerate symmetric or anti-symmetric bilinear forms giving rise to the Grothendieck-Witt group and hermitian $K$-theory. Motivated by direct sum $K$-theory we consider in this article a version of direct sum hermitian $K$-theory. Given a symmetric monoidal category ${\mathcal{C}}$ with duality, e.g. finitely generated projective $R$-modules ${\mathcal{P}}(R)$ over a commutative ring $R$ with the direct sum as symmetric monoidal structure and the usual duality, one can consider the maximal ${\mathbf{E}}_{\infty}$-space in the symmetric monoidal category ${\mathcal{H}}^{\mathrm{nd}}({\mathcal{C}})$ of non-degenerate hermitian objects of ${\mathcal{C}}$ and the spectrum associated to its group completion. This version of hermitian $K$-theory coincides with the common definitions in the case of ${\mathcal{P}}(R)$ (see [30] and [31, Theorem A.1] if $2$ is invertible in $R$ and [20, Theorem A] in general, compare also to [22, 1.6. Theoreme]).
Direct sum $K$-theory is also meaningful in the case of modules over ${\mathbf{E}}_{\infty}$-ring spectra, e.g. the direct sum $K$-theory of finitely generated projective $R$-modules for $R$ a connective ${\mathbf{E}}_{\infty}$-ring spectrum is the connective algebraic $K$-theory of $R$ (see e.g. [24, Theorem 5]). We have a similar behavior in the case of hermitian $K$-theory (see Remark 5.38). In this article we exhibit perfect or finitely generated projective modules over an ${\mathbf{E}}_{\infty}$-ring spectrum as an $\infty$-category with duality together with the relevant symmetric monoidal structures, to which direct sum hermitian $K$-theory can be applied, thereby producing an ${\mathbf{E}}_{\infty}$-ring spectrum.
@@
By Theorem 1.3 the right hand side of injection (1) is the set of equivalence classes of dualities on the $\infty$-category of compact spectra. By Lin’s Theorem [23, Theorem 1.1] there is an isomorphism $\pi_{0}(S^{BC_{2}})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$, where ${\mathbb{Z}}_{2}$ is the uncountable ring of 2-adic numbers. This implies that there are uncountably many inequivalent dualities on the $\infty$-category of compact spectra.
@@
We heavily use symmetric monoidal $\infty$-categories [25, Definition 2.0.0.7.]. We identify symmetric monoidal $\infty$-categories with commutative monoid objects in $\mathsf{Cat}_{\infty}$ via [25, Example 2.4.2.4.]. We denote symmetric monoidal $\infty$-categories in the form $({\mathcal{C}},\otimes)$ and call ${\mathcal{C}}$ the underlying $\infty$-category. We typically drop $\otimes$ from the notation unless there are several symmetric monoidal structures on ${\mathcal{C}}$, which we need to distinguish.
@@
There is a forgetful functor $\mathrm{Calg}({\mathcal{C}})\to{\mathcal{C}}$ that preserves small limits [25, Corollary 3.2.2.5.].
A symmetric monoidal $\infty$-category is cartesian if the tensor unit is final and $\otimes$ is the product. By [25, 2.4.1.5., 2.4.1.7.] every $\infty$-category ${\mathcal{C}}$ that admits finite products carries a unique cartesian symmetric monoidal structure denoted by $({\mathcal{C}},\times)$ and there is a canonical equivalence $\mathrm{Calg}({\mathcal{C}},\times)\simeq\mathrm{Cmon}({\mathcal{C}}).$ For every symmetric monoidal $\infty$-category ${\mathcal{C}}$ and commutative algebra $A$ in ${\mathcal{C}}$ the $\infty$-category ${\mathrm{Mod}}_{A}({\mathcal{C}})$ inherits a symmetric monoidal structure from ${\mathcal{C}}$, the relative tensor product $\otimes_{A}$, such that the free $A$-module functor is symmetric monoidal (see [25, Theorem 4.5.2.1.] or [12, Corollary 7.4.]).
@@
We often use the $\infty$-categorical Grothendieck construction ((un)straightening) [26, Theorem 3.2.0.1.], which classifies functors $\alpha:{\mathrm{S}}\to\mathsf{Cat}_{\infty}$ for any small $\infty$-category ${\mathrm{S}}$ by cocartesian fibrations $\mathrm{X}\to{\mathrm{S}}$ resulting in an equivalence
@@
where the right hand side is the subcategory of cocartesian fibrations to ${\mathrm{S}}$ and functors over ${\mathrm{S}}$ preserving cocartesian lifts. A functor ${\mathrm{S}}\to\mathsf{Cat}_{\infty}$ takes values in spaces if and only if the corresponding cocartesian fibration is a left fibration [26, Proposition 2.4.2.4.]. If ${\mathrm{S}}$ is a space, the cocartesian fibration $\mathrm{X}\to{\mathrm{S}}$ is the functor $\mathrm{colim}(\alpha)\to\mathrm{colim}(*)\simeq{\mathrm{S}}$ induced by the unique map $\alpha\to*$ in $\mathrm{Fun}({\mathrm{S}},\mathsf{Cat}_{\infty})$ by [26, Corollary 3.3.4.3.]. Straightening over $BC_{2}$ gives an equivalence $\mathsf{Cat}_{\infty}[C_{2}]\simeq\mathsf{Cat}_{\infty/BC_{2}}$ that sends an $\infty$-category with $C_{2}$-action ${\mathcal{D}}$ to ${\mathcal{D}}_{hC_{2}}$.
Let ${\mathrm{S}}$ be an $\infty$-category with $C_{2}$-action. A $C_{2}$-equivariant functor $\beta:{\mathrm{S}}\to\mathsf{Cat}_{\infty}$, where $\mathsf{Cat}_{\infty}$ carries the trivial action, is classified by a $C_{2}$-equivariant cocartesian fibration $\mathrm{X}\to{\mathrm{S}}$, refining the non-equivariant (un)straightening construction: the $C_{2}$-equivariant functor $\beta$ corresponds to a functor ${\mathrm{S}}_{hC_{2}}\to\mathsf{Cat}_{\infty}$ classified by a cocartesian fibration $\mathrm{Y}\to{\mathrm{S}}_{hC_{2}}$ (over $BC_{2}$), which by straigtening over $BC_{2}$ is precisely classified by a $C_{2}$-equivariant functor $\mathrm{X}\to{\mathrm{S}}$ whose underlying functor (after forgetting the $C_{2}$-action) is a cocartesian fibration [26, Proposition 2.4.2.11.], which classifies the functor $\beta:{\mathrm{S}}\to{\mathrm{S}}_{hC_{2}}\to\mathsf{Cat}_{\infty}$.
@@ 2.1 $\infty$-categories with duality
The $C_{2}$-action on ${\mathcal{C}}$ is classified via straightening by a functor $\phi:{\mathcal{C}}_{hC_{2}}\to BC_{2}$ by [26, Corollary 3.3.4.3.]. By [26, Corollary 3.3.3.2.] there is a canonical equivalence ${\mathcal{C}}^{hC_{2}}\simeq\mathrm{Fun}_{BC_{2}}(BC_{2},{\mathcal{C}}_{hC_{2}})$ and the forgetful functor ${\mathcal{C}}^{hC_{2}}\to{\mathcal{C}}$ identifies with the functor $\kappa:\mathrm{Fun}_{BC_{2}}(BC_{2},{\mathcal{C}}_{hC_{2}})\to{\mathcal{C}}$ evaluating at the unique object of $BC_{2}$. Hence the object $\widetilde{X\times\tau(X)}\in({\mathcal{C}},\tau)^{hC_{2}}$ with the claimed properties is the right Kan extension relative to $BC_{2}$ of $X:*\to{\mathcal{C}}$ along the functor $*\to BC_{2}$. By [26, Lemma 4.3.2.13.] this right Kan extension exists and sends the unique object of $BC_{2}$ to $X\times\tau(X)$ if the constant functor ×$*\times_{BC_{2}}*\simeq C_{2}\to{\mathcal{C}}_{hC_{2}}$ with value $X$ admits a $\phi$-limit. By [26, Proposition 4.3.1.9.] this $\phi$-limit is the limit of the functor $C_{2}\to{\mathcal{C}}$ corresponding to the pair $(X,\tau(X))$, which is $X\times\tau(X).$
@@ 2.2 Bimonoidal $\infty$-categories with duality
By [25, Corollary 4.8.1.4, Proposition 4.8.1.17.] there is a closed symmetric monoidal structure on $\mathsf{Pr}^{L}$, where the internal hom of ${\mathcal{C}},{\mathcal{D}}\in\mathsf{Pr}^{L}$ is the $\infty$-category $\mathrm{Fun}^{L}({\mathcal{C}},{\mathcal{D}})$ of left adjoint functors ${\mathcal{C}}\to{\mathcal{D}}$.
@@
An object $\mathrm{X}\in\mathrm{Rig}({\mathcal{C}})$ roughly consists of an object $\mathrm{X}\in{\mathcal{C}}$ equipped with two coherently unital, associative and commutative operations $+,\cdot:\mathrm{X}\times\mathrm{X}\to\mathrm{X}$ such that $\cdot$ coherently distributes over +.+.+ .
@@
The $\infty$-category $\mathsf{Cat}_{\infty}^{\prod}$ admits finite products (in fact all small limits) that are preserved by the inclusion $\mathsf{Cat}_{\infty}^{\prod}\subset\mathsf{Cat}_{\infty}$. Since every finite products preserving functor preserves the final object, the $\infty$-category $\mathsf{Cat}_{\infty}^{\prod}$ admits a zero object. For every ${\mathcal{C}}\in\mathsf{Cat}_{\infty}^{\prod}$ let $\mu_{\mathcal{C}}:{\mathcal{C}}\times{\mathcal{C}}\to{\mathcal{C}}$ the functor assigning the two-fold product. By [25, Proposition 2.4.3.19.] the statement follows from the fact that the functor ${\mathcal{C}}\times 0\to{\mathcal{C}}\times{\mathcal{C}}\xrightarrow{\mu_{\mathcal{C}}}{\mathcal{C}}$ is the identity, the functor $({\mathcal{C}}\times{\mathcal{C}})\times({\mathcal{D}}\times{\mathcal{D}})\simeq({\mathcal{C}}\times{\mathcal{D}})\times({\mathcal{C}}\times{\mathcal{D}})\xrightarrow{\mu_{{\mathcal{C}}\times{\mathcal{D}}}}{\mathcal{C}}\times{\mathcal{D}}$ is the functor $\mu_{\mathcal{C}}\times\mu_{\mathcal{D}}$ and for every map $\mathsf{f}:{\mathcal{C}}\to{\mathcal{D}}$ in $\mathsf{Cat}_{\infty}^{\prod}$ we have $\mathsf{f}\circ\mu_{\mathcal{C}}=\mu_{\mathcal{D}}\times(\mathsf{f}\times\mathsf{f}).$
@@
This follows since $\mathsf{Cat}_{\infty}^{\mathrm{preadd}}\subset\mathsf{Cat}_{\infty}^{\prod}$ is closed under finite products.
@@
By [25, Corollary 4.8.1.4] the $\infty$-category $\mathsf{Cat}_{\infty}^{\prod}$ carries a presentably symmetric monoidal structure such that the inclusion $\mathsf{Cat}_{\infty}^{\prod}\subset\mathsf{Cat}_{\infty}$ admits a symmetric monoidal left adjoint. Applying the lax symmetric monoidal localization functor $\mathrm{Cmon}:\mathsf{Pr}^{L}\to\mathsf{Pr}^{L}$ we find that the embedding (2) admits a symmetric monoidal left adjoint and so refines to a lax symmetric monoidal embedding
@@
Let $\mathsf{Cat}_{\infty}^{\coprod}\subset\mathsf{Cat}_{\infty}$ be the subcategory of small $\infty$-categories having finite coproducts and functors preserving finite coproducts. Taking the opposite $\infty$-category restricts to an equivalence $\mathsf{Cat}_{\infty}^{\prod}\simeq\mathsf{Cat}_{\infty}^{\coprod}$, which further restricts to an autoequivalence of $\mathsf{Cat}_{\infty}^{\mathrm{preadd}}.$ Consequently, it is enough to prove that the embedding $\mathsf{Cat}_{\infty}^{\mathrm{preadd}}\subset\mathsf{Cat}_{\infty}^{\coprod}$ admits a symmetric monoidal left adjoint, which is notationally more convenient. For any small $\infty$-category ${\mathcal{C}}$ that admits finite coproducts let ${\mathcal{C}}^{\prime}$ be the essential image of the composition ${\mathcal{C}}\to\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})\to\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))\simeq\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$ of the Yoneda-embedding and the free functor. Since the Yoneda-embedding ${\mathcal{C}}\to\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ preserves finite coproducts [26, Proposition 5.5.8.10.], the full subcategory ${\mathcal{C}}^{\prime}$ is closed in the preadditive $\infty$-category $\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))$ under finite coproducts and therefore is itself preadditive. We will prove that the induced functor ${\mathcal{C}}\to{\mathcal{C}}^{\prime}$ induces an equivalence $\kappa_{\mathcal{D}}:\mathrm{Fun}^{\coprod}({\mathcal{C}}^{\prime},{\mathcal{D}})\to\mathrm{Fun}^{\coprod}({\mathcal{C}},{\mathcal{D}}).$ The functor $\kappa_{\mathcal{D}}$ is the pullback of the functor $\kappa_{\mathrm{Fun}^{\prod}({\mathcal{D}}^{\mathrm{op}},{\mathcal{S}})}$ along the Yoneda-embedding ${\mathcal{D}}\subset\mathrm{Fun}^{\prod}({\mathcal{D}}^{\mathrm{op}},{\mathcal{S}})\simeq\mathrm{Fun}({\mathcal{D}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$. Since $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ is preadditive and presentable [26, Proposition 5.5.8.10.], we can therefore assume that ${\mathcal{D}}$ is a presentable preadditive $\infty$-category when proving that $\kappa_{\mathcal{D}}$ is an equivalence.
@@
The $\infty$-category $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ is generated by ${\mathcal{C}}$ under small sifted colimits [26, Lemma 5.5.8.14.]. Since $\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))$ is generated by $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ under small sifted colimits, $\mathrm{Cmon}(\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}}))$ is generated by ${\mathcal{C}}^{\prime}$ under small sifted colimits. Every object of ${\mathcal{C}}$ is compact and projective in $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},{\mathcal{S}})$ by [26, Proposition 5.5.8.10.4.] so that every object of ${\mathcal{C}}^{\prime}$ is compact and projective in $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$. By [26, Proposition 5.5.8.22.] this guarantees that the embedding ${\mathcal{C}}^{\prime}\subset\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}}))$ induces an equivalence $\mathrm{Fun}^{\prod}({\mathcal{C}}^{\prime\mathrm{op}},{\mathcal{S}})\simeq\mathrm{Fun}^{\prod}({\mathcal{C}}^{\mathrm{op}},\mathrm{Cmon}({\mathcal{S}})).$
@@
For every small $\infty$-category ${\mathcal{B}}$ that admits finite coproducts the induced functor $\mathrm{Fun}^{\mathrm{L}}(\mathrm{Fun}^{\prod}({\mathcal{B}}^{\mathrm{op}},{\mathcal{S}}),{\mathcal{D}})\to\mathrm{Fun}^{\coprod}({\mathcal{B}},{\mathcal{D}})$ is an equivalence [26, Proposition 5.5.8.15.]. So in the following commutative square the vertical functors are equivalences:
@@
whose fully faithful right adjoint identifies preadditive symmetric monoidal $\infty$-categories with duality $({\mathcal{C}},\otimes,D)$ with preadditive bimonoidal $\infty$-categories with duality $({\mathcal{C}},\otimes,\boxtimes,D)$, where $\boxtimes$=$\oplus$ = ⊕ is the sum.